OpenStudy (anonymous):

Solve the following system of equations. Express your answer as an ordered pair in the format open parentheses a comma b close parentheses, with no spaces between the numbers or symbols. 3x+6y=3 2x-10y=44

pooja195 (pooja195):

Hi welcome to openstudy :) So the problem gives use 2 equations 3x+6y=3 <---equation 1 2x-10y=44 <---equation 2 let's isolate x in the first equation so that we can solve by substitution Can you please isolate the x? :) 3x+6y=3 Make sure you show the steps

OpenStudy (anonymous):

3x+6y=3 -6y -6y 3x=3-6y -------- 3 3 3 x=1-2y is that correct?

pooja195 (pooja195):

Good personally i would right it as -2y+1 means the exact same thing x=-2y+1 <---equation 1 2x-10y=44 <---equation 2 Now we can solve by substitution equation 1 into equation 2 2(-2y+1)-10y=44 solve for y

OpenStudy (anonymous):

2(-2y+1)-10y=44 -47+2-10y=44 +10y -2 -4y=44+10y-2 -------------- -4 -4 -4 -4 y=-11-2.5y-.5

pooja195 (pooja195):

Im not sure what you did.... we get a hole number for the value of x

OpenStudy (anonymous):

what do u mean

OpenStudy (anonymous):

what was i supposed to do?

pooja195 (pooja195):

2(-2y+1)-10y=44 start by distribtuing the 2 2(-2y+1)

OpenStudy (anonymous):

i did that which gave me -4y+2

pooja195 (pooja195):

-4y+2-10y=44 combine like terms -4-10=?

OpenStudy (anonymous):

-14

pooja195 (pooja195):

good −14y+2=44 subtract 2 from both sides what are you left with?

OpenStudy (anonymous):

-16y=42???

OpenStudy (anonymous):

wait or -14y=42

pooja195 (pooja195):

Yes didive both sides by -14 y=?

OpenStudy (anonymous):

y=-3.14 thats pi

pooja195 (pooja195):

42/-14=?

OpenStudy (anonymous):

close enough lol

OpenStudy (anonymous):

-3 my bad

pooja195 (pooja195):

good so y=-3 x=-2y+1 Plug -3 in place of y x=-2(-3)+1

pooja195 (pooja195):

Solve

OpenStudy (anonymous):

x=7

OpenStudy (anonymous):

so my answer is (7,-3)

pooja195 (pooja195):

Good job :) so we have x=7 y=-3 Correct

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

THANK YOU SO MUCH

pooja195 (pooja195):

yw ^_^